# Spin-orbit misalignment and residual eccentricity are evidence that neutron star-black hole mergers form through triple star evolution

### v1

@@ 3 Results
The properties (i) – (v) naturally result in a large fraction of tertiary-driven NSBH mergers with residual eccentricity and spin-orbit misalignment of the BH, as shown in Fig. 9. The left panel shows that a high fraction ($\sim 60\,\%$) tertiary-driven NSBH mergers retain a large residual eccentricity ($e_{20}\gtrsim 0.1$) at $f_{\rm GW}=20\,\rm Hz$. The detectable range of eccentricities with the current network of ground-based GW detectors is about $0.07\lesssim e_{20}\lesssim 0.4$, where below eccentricities are too small to be confidently distinguishable from circular orbits (M. E. Lower et al., 2018; A. H. Nitz et al., 2020) and above (yellow patches) the GW signal becomes burst-like and is not well-modelled by some of current eccentric waveform approximants (e.g., A. H. Nitz et al., 2020; M. de Lluc Planas et al., 2025). Moreover, highly eccentric events risk being missed by the current searches for circular events by the LVK and would therefore not recognised by G. Morras et al. (2025) (although using approximants sensitive to higher eccentricities) who only reanalysed events which were identified by the LVK in the first place. Restricting the merger sample to the subset with $e_{20}\lesssim 0.4$, we find a fraction of $\sim 50\,\%$ with $e_{20}\gtrsim 0.07$ (red line and patches). Obtaining an eccentricity at least as large as $e_{20}\geq 0.145$ as inferred by G. Morras et al. (2025) for GW200105 has a probability of $\sim 25\,\%$. Therefore, detecting such eccentric event out of up to six NSBH mergers detected so far would be a very plausible outcome if triples are the dominant formation channel. Meanwhile, we expect a fraction $\sim 96\,\%$ to have eccentricities $e_{20}\gtrsim 10^{-3}$ which would be potentially detectable with the Einstein Telescope and Cosmic Explorer (P. Saini, 2024). In contrast, any event with detectable eccentricity would be wholly inconsistent with mergers originating from isolated NSBH binaries (blue patches and teal line) and would rule them out as a significant formation channel.
@@
As shown on the right panel of Fig. 9, the distribution of the final BH spin-orbit angles of tertiary-driven NSBH mergers extends to large misalignment values ($\cos\theta_{\rm BH}\approx-1$) but peaks near alignment ($\cos\theta_{\rm BH}\approx+1$). This distribution can be understood by the two cases described in (ii) and exemplified in Figs. 7 and 8, where mergers with close spin-orbit alignment result from NSBH triples that could not have completed a full ZKL cycle before the mergers take place, whereas large misalignments are the result of large-amplitude oscillations of the orbital orientation across several ZKL cycles in the non-adiabatic regime. Obtaining retrograde spins with $\cos\theta_{\rm BH}<0$ as found for GW200115 (R. Abbott et al., 2021) and GW230529 (A. G. Abac et al., 2024) is the case for $\sim 14\,\%$ of the tertiary-driven mergers, which would be in agreement with $\sim 1$ – $2$ such events out of $\lesssim 6$ detected NSBH candidates. In our models this is $4$ – $5$ times more likely than for isolated binary mergers (although see G. Fragione et al., 2021).
@@
Finally, the fact that large octupole effects (see (i)) enable mergers across a wide range of mutual inclinations (Fig. 5) leads to a large fraction and rate of NSBH mergers which agrees with the current estimate $\mathcal{R}_{\rm LVK}=55^{+127}_{-47}\,\rm Gpc^{-3}\,yr^{-1}$ from GW detections (A. G. Abac et al., 2024). We follow G. Fragione & A. Loeb (2019a) and calculate the NSBH merger rate by assuming a local star formation rate of $0.025\,\rm M_{\odot}\,Gpc^{-3}\,yr^{-1}$ so that the number of primary stars formed per unit mass is given by (M. S. Bothwell et al., 2011)
@@
where we assume that a fraction $f_{3}=0.75$ of massive stars are in hierarchical triples on the ZAMS (M. Moe & R. Di Stefano, 2017) and determine the fraction $f_{\rm merge}=N_{\rm merge}/N_{\rm tot}$ from the number $N_{\rm merge}$ of tertiary-driven NSBH mergers formed out of $N_{\rm tot}$ stellar triples obtained in our simulations. Depending on the kick model we obtain
$\mathcal{R}_{\text{NSBH}}$ =$=\begin{cases}5.46&\text{Maxwellian\_265\_30}\\ 22.87&\text{Maxwellian\_100\_10}\\ 219.24&\text{Zero}\\ 6.30&\text{Igoshev}\\ 22.81&\text{Valli}\\ 2.70&\text{Disberg}\\ 1.31&\text{Bray}\\ \end{cases}\quad\text{Gpc}^{-3}~{}\text{yr}^{-1},$
@@
$\mathcal{R}_{\text{NSBH}}^{e_{20}<0.4}$ =$=\begin{cases}4.02&\text{Maxwellian\_265\_30}\\ 15.91&\text{Maxwellian\_100\_10}\\ 133.67&\text{Zero}\\ 4.29&\text{Igoshev}\\ 15.88&\text{Valli}\\ 2.05&\text{Disberg}\\ 0.89&\text{Bray}\\ \end{cases}\quad\text{Gpc}^{-3}~{}\text{yr}^{-1}.$
In addition, some of the BHOB inner binaries may undergo CE evolution and form NSBH mergers essentially in isolation, regardless of the tertiary. Their contribution would increase the total rate and dilute somewhat the fraction of eccentric events. However, the question which systems survive CE evolution with post-CE orbits that are close enough to produce GW events is hugely uncertain. This is not only because of the general complexity of the CE phase (N. Ivanova et al., 2020), but also due to the poorly-understood interplay between the inner binary and the rich circumbinary matter ejected during the dynamical phase of CE, which may strongly influence the final orbit (P. A. Ondratschek et al., 2022; D. Gagnier & O. Pejcha, 2023; S. Tuna & B. D. Metzger, 2023; D. Wei et al., 2024; M. Vetter et al., 2024, 2025). The specific case of NSBHs might be particularly challenging for the CE channel if the dynamical in-spiral is followed by a slower mass exchange during the self-regulated phase (N. Ivanova, 2011; T. Fragos et al., 2019; A. Vigna-Gómez et al., 2022; R. Hirai & I. Mandel, 2022), which for unequal mass ratios of NSBHs is likely to widen the final orbit, easily to the point when the system no longer merges as a GW source (M. M. Moreno et al., 2022; D. Wei et al., 2024). The contribution of the CE channel to the NSBH merger population, if any, will be constrained by future GW events.
@@ 4 Conclusion
Deciphering the origin of the compact object mergers observed through GW radiation is one of the major challenges in GW astronomy. The formation of NSBH mergers is particularly intriguing as the natal kick associated with the NS formation and their low mass ratio pose a challenge to proposed mechanisms, e.g., few-body scatterings in dense star clusters (see Sec. 1). Meanwhile, formation from the evolution of isolated massive binary stars would be wholly inconsistent with the detection of eccentric NSBH mergers and is challenged by the detection of events with large BH spin-orbit misalignment, as recently inferred for GW200105 (G. Morras et al., 2025; M. de Lluc Planas et al., 2025) and GW200115 and GW230529 (R. Abbott et al., 2021; A. G. Abac et al., 2024), respectively.
@@
Motivated by the fact that most massive stars reside in hierarchical triple systems (M. Moe & R. Di Stefano, 2017), we investigated their evolution into NSBH mergers driven by gravitational perturbations from the tertiary companion via the ZKL mechanism. Among the various stages of stellar evolution, the second SN—which forms the NS—is the most disruptive event to the stability of the triple system. As a result, only compact NSBH triples tend to survive, characterised by properties (i) – (v). Thus, we find the NSBH mergers in triples to occur at a rate $\mathcal{R}_{\rm NSBH}\simeq 1$ – $23\,\rm Gpc^{-3}\,yr^{-1}$ ($\mathcal{R}^{e_{20}<0.4}_{\rm NSBH}\simeq 1$ – $16\,\rm Gpc^{-3}\,yr^{-1}$) and we obtain a considerable fraction of a few $10\,\%$ of residual eccentricity and large BH spin-orbit misalignment (Fig. 9). This is in excellent agreement with the rate of NSBH mergers inferred by the LVK as well as the number and the fraction of peculiar GW events showing tentative evidence for eccentricity and misalignment. This will be further constrained by NSBH detections in the upcoming GW observing runs. Further evidence for the triple channel may come from the growing sample of BHOB systems, possibly with tertiary companions (Klencki & Stegmann, in prep.), that is set to substantially increase thanks to large-scale astrometric (GAIA), spectroscopic (e.g. LAMOST, SDSS), and microlensing surveys (e.g., Vera Rubin, N. S. Abrams et al., 2025; T. Shenar et al., 2024; K. El-Badry, 2024).
@@
If GW200105 is indeed eccentric or similar events are confidently detected in the future, our work suggests that hierarchical field triples are the dominant formation channel of NSBH mergers. Nonetheless, close binary star evolution on the inner orbit is a crucial aspect in determining the properties of the resulting NSBH mergers. On the one hand, a close separation ($a_{1}\lesssim\mathcal{O}(10)\,\rm AU$) of the inner binary is needed for a large number to survive the natal kick of the NS. Not including them explains why previous work by G. Fragione & A. Loeb (2019a) only found a high rate of NSBH mergers from triples if zero kicks were assumed. On the other hand, close binaries interact through mass transfer which critically affects the properties of the resulting compact remnants. We leave a detailed study investigating the implications of different assumptions about binary interaction for future work.
@@
If binary black holes would predominantly merge in hierarchical triples too, we speculate it may provide a natural explanation for the seemingly smaller fraction of a few events out of about a hundred that have tentative evidence for eccentricity (I. Romero-Shaw et al., 2022; N. Gupte et al., 2024) and negative $\chi_{\rm eff}$ (R. Abbott et al., 2023). On the one hand, the mass-loss and natal kick associated with BH formation is generally expected to be much smaller than for NSs allowing much wider triple orbits to survive (e.g., F. Antonini et al., 2017; C. L. Rodriguez & F. Antonini, 2018). On the other hand, the mass ratio distribution of binary black holes is generally closer to one compared to NSBHs, which would diminish the strength of octupole effects and narrow the parameter window for tertiary-driven mergers. As a result, we would expect the fraction of eccentric binary black hole mergers and with spin-orbit misalignment to be much lower than for NSBHs if both systems form in hierarchical triples.
@@
We thank Patricia Schmidt, Fabio Antonini, Yossef Zenati, Alejandro Vigna-Gómez, Ruggero Valli, and Martyna Chruślińska for useful discussions and input. {contribution} Both authors contributed equally.

### v2

@@ III Results
The properties (i) – (v) naturally result in a large fraction of tertiary-driven NSBH mergers with residual eccentricity and spin-orbit misalignment of the BH, as shown in Fig. 7, in the Maxwellian_100_10 model. The left panel shows that a high fraction ($\sim 60\,\%$) tertiary-driven NSBH mergers retain a large residual eccentricity ($e_{20}\gtrsim 0.1$) at $f_{\rm GW}=20\,\rm Hz$. The detectable range of eccentricities with the current network of ground-based GW detectors is about $0.07\lesssim e_{20}\lesssim 0.4$, where below eccentricities are too small to be confidently distinguishable from circular orbits (cf., Lower et al., 2018; Nitz et al., 2020, on eccentric binary black hole mergers at $f_{\rm GW}=10\,\rm Hz$) and above (yellow patches) the GW signal becomes burst-like and is not well-modelled by some of current eccentric waveform approximants (e.g., Nitz et al., 2020; de Lluc Planas et al., 2025). Moreover, highly eccentric events risk being missed by the current searches for circular events by the LVK and would therefore not recognised by Morras et al. (2025) (although using approximants sensitive to higher eccentricities) who only reanalysed events which were identified by the LVK in the first place. Restricting the merger sample to the subset with $e_{20}\lesssim 0.4$, we find a fraction of $\sim 50\,\%$ with $e_{20}\gtrsim 0.07$ (red line and patches). Obtaining an eccentricity at least as large as $e_{20}\geq 0.145$ as inferred by Morras et al. (2025) for GW200105 has a probability of $\sim 25\,\%$. Therefore, detecting such eccentric event out of up to seven NSBH mergers detected so far would be a very plausible outcome if triples are the dominant formation channel. Meanwhile, we expect a fraction $\sim 96\,\%$ to have eccentricities $e_{20}\gtrsim 10^{-3}$ which would be potentially detectable with the Einstein Telescope and Cosmic Explorer (Saini, 2024). In contrast, any event with detectable eccentricity would be wholly inconsistent with mergers originating from isolated NSBH binaries (blue patches and teal line) and would rule them out as a significant formation channel.
The above findings do not largely change when switching from the secular to the direct $N$-body integration, shown in dashed lines in the left panel of Fig. 7. We find that the fraction of mergers the fraction of eccentric mergers drops by a factor of $\sim 2$ compared to the secular integration. Moreover, the total number of mergers is found to be smaller by a factor of $\sim 5$. Individual systems often merge with one integration technique but not with the other, due to non-secular torques in the direct $N$-body simulation. However, the overall population statistics remain somewhat similar, although we do not find an increased total number of mergers and eccentric ones, as found in previous studies (e.g., Antonini et al., 2014).
@@
As shown on the right panel of Fig. 7, the distribution of the final BH spin-orbit angles of tertiary-driven NSBH mergers extends to large misalignment values ($\cos\theta_{\rm BH}\approx-1$) but peaks near alignment ($\cos\theta_{\rm BH}\approx+1$). This distribution can be understood by the two cases described in (ii) and exemplified in Figs. 6 and 10, where mergers with close spin-orbit alignment result from NSBH triples that could not have completed a full ZKL cycle before the mergers take place, whereas large misalignments are the result of large-amplitude oscillations of the orbital orientation across several ZKL cycles in the non-adiabatic regime. Obtaining retrograde spins with $\cos\theta_{\rm BH}<0$ as found for GW200115 (Abbott et al., 2021) and GW230529 (Abac et al., 2024) is the case for $\sim 14\,\%$ of the tertiary-driven mergers, which would be in agreement with $\sim 1$ – $2$ such events out of $\lesssim 6$ detected NSBH candidates. In our models this is $4$ – $5$ times more likely than for isolated binary mergers (although see Fragione et al., 2021).
@@
Finally, the fact that large octupole effects (see (i)) enable mergers across a wide range of mutual inclinations (Fig. 4) leads to a large fraction and rate of NSBH mergers which agrees with the current estimate $\mathcal{R}_{\rm LVK}=9.1$ – $84\,\rm Gpc^{-3}\,yr^{-1}$ from GW detections (The LIGO Scientific Collaboration et al., 2025a). We follow Fragione and Loeb (2019b) and calculate the NSBH merger rate by assuming a local star formation rate of $0.025\,\rm M_{\odot}\,Gpc^{-3}\,yr^{-1}$ so that the number of primary stars formed per unit mass is given by (Bothwell et al., 2011)
@@
where we assume that a fraction $f_{3}=0.75$ of massive stars are in hierarchical triples on the ZAMS (Moe and Di Stefano, 2017) and determine the fraction $f_{\rm merge}=N_{\rm merge}/N_{\rm tot}$ from the number $N_{\rm merge}$ of tertiary-driven NSBH mergers formed out of $N_{\rm tot}$ stellar triples obtained in our simulations. Depending on the kick model we obtain
$\mathcal{R}_{\text{NSBH}}$ $=\begin{cases}5.46&\text{Maxwellian\_265\_30}\\ 22.87&\text{Maxwellian\_100\_10}\\ 219.24&\text{Zero}\\ 6.30&\text{Igoshev}\\ 22.81&\text{Valli}\\ 2.70&\text{Disberg}\\ 1.31&\text{Bray}\\ \end{cases}\quad\text{Gpc}^{-3}~\text{yr}^{-1},$
@@
$\mathcal{R}_{\text{NSBH}}^{e_{20}<0.4}$ $=\begin{cases}4.02&\text{Maxwellian\_265\_30}\\ 15.91&\text{Maxwellian\_100\_10}\\ 133.67&\text{Zero}\\ 4.29&\text{Igoshev}\\ 15.88&\text{Valli}\\ 2.05&\text{Disberg}\\ 0.89&\text{Bray}\\ \end{cases}\quad\text{Gpc}^{-3}~\text{yr}^{-1}.$
The above range of NSBH merger rates illustrates the extent to which this channel is affected by natal kicks, but it should not be taken as the overall rate uncertainty of the triple scenario. While natal kicks are a key factor, other poorly constrained processes also influence the survival of massive-star triples on their evolutionary path to forming NSBH triples. In particular, these include: (i) the stability of mass transfer in the inner binary, (ii) the stability of the triple is response to inner binary widening, (iii) and the criteria for the formation of a BH or a NS. In the model of Schürmann and Langer (2024) adopted here, the stability of mass transfer is directly related to the accretion efficiency ($\beta$): the more a star accretes, the more likely it is to rapidly expand in response and trigger instability. For our fiducial model, we adopted $\beta=0.1$. motivated by the broad expectation for rotationally-limit accretion (Langer et al., 2003). Observations of post-interaction binaries suggest that a broader range of $\beta$ is possible (Lechien et al., 2025). We find that when instead assuming $\beta=0.7$ ($0.0$), the number of NSBH triples in our population model decreases (increases) by a factor of $\sim 4$. To gauge the effect of (ii), we explore a variation with a ’Jeans’ mode for the ejection of the non-accreted matter, leading to more substantial inner orbit widening. We find that the number of NSBH triples decreases by only $\sim$10%. Finally, to test the effect of (iii), we replace the prescription for BH/NS formation from Woosley et al. (2020) with that of Fryer et al. (2012, their Delayed variant) and Maltsev et al. (2025, their Model A variant). We find that the predicted number of NSBH triples decreases by a factor $\sim 2$ and $\sim 3$, respectively. We therefore heuristically estimate that the NSBH rates given above for different natal kick variations are each subject to further uncertainty by a factor of $\sim 10$. Finally, our rates are likely subject to further uncertainty due to the poorly-constrained initial parameters of massive hierarchical triples. However, thanks in part to the significant octupole terms of the ZKL effect, our NSBH mergers are not restricted to any narrow parameter range of the triple orbits (e.g. no fine-tuning of inclinations). The most constrained parameter is the ratio $a_{2}/a_{1}$, ranging from $\sim$5 to $\sim$500 (Fig. 2). While there are no strong constraints on the $a_{2}/a_{1}$ distribution for massive triples, among low-mass triples the $a_{2}/a_{1}$ was found to peak around a favourable value $\sim 10$ (Shariat et al., 2025).
In addition, some of the BHOB inner binaries may undergo CE evolution and form NSBH mergers essentially in isolation, regardless of the tertiary. Their contribution would increase the total rate and dilute somewhat the fraction of eccentric events. However, the question which systems survive CE evolution with post-CE orbits that are close enough to produce GW events is hugely uncertain. This is not only because of the general complexity of the CE phase (Ivanova et al., 2020), but also due to the poorly-understood interplay between the inner binary and the rich circumbinary matter ejected during the dynamical phase of CE, which may strongly influence the final orbit (Ondratschek et al., 2022; Gagnier and Pejcha, 2023; Tuna and Metzger, 2023; Wei et al., 2024; Vetter et al., 2024, 2025). The specific case of NSBHs might be particularly challenging for the CE channel if the dynamical in-spiral is followed by a slower mass exchange during the self-regulated phase (Ivanova, 2011; Fragos et al., 2019; Vigna-Gómez et al., 2022; Hirai and Mandel, 2022), which for unequal mass ratios of NSBHs is likely to widen the final orbit, easily to the point when the system no longer merges as a GW source (Moreno et al., 2022; Wei et al., 2024). The contribution of the CE channel to the NSBH merger population, if any, will be constrained by future GW events.
@@ IV Conclusion
Deciphering the origin of the compact object mergers observed through GW radiation is one of the major challenges in GW astronomy. The formation of NSBH mergers is particularly intriguing as the natal kick associated with the NS formation and their low mass ratio pose a challenge to proposed mechanisms, e.g., few-body scatterings in dense star clusters (see Sec. I). Meanwhile, formation from the evolution of isolated massive binary stars would be wholly inconsistent with the detection of eccentric NSBH mergers and is challenged by the detection of events with large BH spin-orbit misalignment, as recently inferred for GW200105 (Fei and Yang, 2024; Morras et al., 2025; de Lluc Planas et al., 2025; Kacanja et al., 2025) and GW200115 and GW230529 (Abbott et al., 2021; Abac et al., 2024), respectively.
@@
Motivated by the fact that most massive stars reside in hierarchical triple systems (Moe and Di Stefano, 2017), we investigated their evolution into NSBH mergers driven by gravitational perturbations from the tertiary companion via the ZKL mechanism. Among the various stages of stellar evolution, the second SN—which forms the NS—is the most disruptive event to the stability of the triple system. As a result, only compact NSBH triples tend to survive, characterised by properties (i) – (v). Thus, we find the NSBH mergers in triples to occur at a rate $\mathcal{R}_{\rm NSBH}\simeq 1$ – $23\,\rm Gpc^{-3}\,yr^{-1}$ ($\mathcal{R}^{e_{20}<0.4}_{\rm NSBH}\simeq 1$ – $16\,\rm Gpc^{-3}\,yr^{-1}$) and we obtain a considerable fraction of a few $10\,\%$ of residual eccentricity and large BH spin-orbit misalignment (Fig. 7). This is in excellent agreement with the rate of NSBH mergers inferred by the LVK as well as the number and the fraction of peculiar GW events showing tentative evidence for eccentricity and misalignment. This will be further constrained by NSBH detections in the upcoming GW observing runs. Further evidence for the triple channel may come from the growing sample of BHOB systems, possibly with tertiary companions (Klencki & Stegmann, in prep.), that is set to substantially increase thanks to large-scale astrometric (GAIA), spectroscopic (e.g. LAMOST, SDSS), and microlensing surveys (e.g., Vera Rubin, Abrams et al., 2025; Shenar et al., 2024; El-Badry, 2024).
@@
If GW200105 is indeed eccentric or similar events are confidently detected in the future, our work suggests that hierarchical field triples are the dominant formation channel of NSBH mergers. Nonetheless, close binary star evolution on the inner orbit is a crucial aspect in determining the properties of the resulting NSBH mergers. On the one hand, a close separation ($a_{1}\lesssim\mathcal{O}(10)\,\rm AU$) of the inner binary is needed for a large number to survive the natal kick of the NS. Not including them explains why previous work by Fragione and Loeb (2019b) only found a high rate $\mathcal{R}_{\rm NSBH}\simeq 19\,{\rm Gpc}^{-3}\,{\rm yr}^{-1}$ of NSBH mergers from triples if zero kicks were assumed and $\mathcal{R}_{\rm NSBH}\lesssim\mathcal{O}(10^{-2})\,{\rm Gpc}^{-3}\,{\rm yr}^{-1}$ if kicks taken into account. Meanwhile, previous work by Hamers and Thompson (2019) did find a high rate of NSBH mergers up to $\mathcal{R}_{\rm NSBH}\lesssim\mathcal{O}(10^{2})\,{\rm Gpc}^{-3}\,{\rm yr}^{-1}$, but since they did not fully account for three-body dynamics they could not report on the fraction of mergers with eccentricity or spin-orbit misalignment. On the other hand, close binaries interact through mass transfer which critically affects the properties of the resulting compact remnants. We leave a detailed study investigating the implications of different assumptions about binary interaction for future work.
@@
If binary black holes would predominantly merge in hierarchical triples too, we speculate it may provide a natural explanation for the seemingly smaller fraction of a few events out of about a hundred that have tentative evidence for eccentricity (Romero-Shaw et al., 2022; Gupte et al., 2024; Romero-Shaw et al., 2025) and negative $\chi_{\rm eff}$ (Abbott et al., 2023). On the one hand, the mass-loss and natal kick associated with BH formation is generally expected to be much smaller than for NSs allowing much wider triple orbits to survive (e.g., Antonini et al., 2017; Rodriguez and Antonini, 2018). On the other hand, the mass ratio distribution of binary black holes is generally closer to one compared to NSBHs, which would diminish the strength of octupole effects and narrow the parameter window for tertiary-driven mergers. As a result, we would expect the fraction of eccentric binary black hole mergers and with spin-orbit misalignment to be much lower than for NSBHs if both systems form in hierarchical triples.
@@
We thank Patricia Schmidt, Fabio Antonini, Yossef Zenati, Alejandro Vigna-Gómez, Ruggero Valli, Martyna Chruślińska, and Isobel Romero-Shaw for useful discussions and input.
{contribution}
Both authors contributed equally.
Appendix A Supplementary figures
Figure 10: Same as Fig. 6 for an example system that merges before fully completing a full ZKL cycle. As a result the inner orbital angular momentum could only slightly deviate from the BH spin direction. The initial parameters are $m_{1}=7.30\,\rm M_{\odot}$, $m_{2}=1.50\,\rm M_{\odot}$, $m_{3}=0.13\,\rm M_{\odot}$, $e_{1}=0.58$, $a_{1}=2.22\,\rm AU$, $e_{2}=0.63$, $a_{2}=58.22\,\rm AU$, $\mathbf{\hat{e}}_{1}=(-0.92,0.17,0.35)^{\rm T}$, $\mathbf{\hat{j}}_{1}=(-0.33,0.11,-0.94)^{\rm T}$, $\mathbf{\hat{e}}_{2}=(-0.51,0.55,-0.66)^{\rm T}$, $\mathbf{\hat{j}}_{2}=(-0.19,0.68,0.71)^{\rm T}$, and $\mathbf{\hat{S}}_{\rm BH}=(-0.33,0.17,-0.93)^{\rm T}$.